OpenStudy (anonymous):

Darlene kicks a soccer ball off the ground and in the air, with an initial velocity of 34 feet per second. Using the formula H(t) = −16t2 + vt + s, what is the maximum height the soccer ball reaches? 17.7 feet 18.1 feet 19.3 feet 20.2 feet

OpenStudy (freckles):

If you find the vertex of the parabola you can find the max height.

OpenStudy (freckles):

\[H(t)=-16t^2+v_0t+h_0 \\ \text{ where } v_0 \text{ is initial velocity } \\ \text{ and } h_0 \text{ is the initial height } \\ \text{ you are given } v_0=34 \frac{\text{ ft}}{\text{ sec}} \\ \text{ and the inital height was given as zero since that is where the ball started at } \\ \text{ you know on the ground }\]

OpenStudy (freckles):

so plug in the numbers

OpenStudy (freckles):

and then write into the vertex form let me know if you need any help

OpenStudy (anonymous):

So the answer is C correct or did i do the math wrong.

OpenStudy (freckles):

not sure what did you do?

OpenStudy (freckles):

\[ax^2+bx+c \\ a(x^2+\frac{b}{a}x)+c \\ a(x^2+\frac{b}{a}x+(\frac{b}{2a})^2)+c-a(\frac{b}{2a})^2 \\ a(x+\frac{b}{2a})^2+c-a(\frac{b}{2a})^2 \\ \text{ vertex of } ax^2+bx+c \text{ is } (-\frac{b}{2a},c-a (\frac{b}{2a})^2) \\ \text{ the \max is given by the y-coordinate }\]

OpenStudy (anonymous):

The answer is 18.1 :}